Jefferson Lab HALL A Logbook for October 01 - All EXCEPT Automatic Entries

  • Main INDEX

    Wednesday

  • 75521 10/31/01 23:48 johna Change to physics kumac.
  • 75508 10/31/01 17:52 johna Detector kumac changes.
  • 75503 10/31/01 16:10 jpchen 10/31 day shift summary
  • 75501 10/31/01 15:54 LeRose OTR non-functional
  • 75498 10/31/01 15:05 LeRose Blank displays, the right answer
  • 75497 10/31/01 14:59 jpchen Hall A general tools screen
  • 75494 10/31/01 14:10 LeRose Beamline display blank
  • 75491 10/31/01 12:47 R. Michaels scaler measure of EDT nonsense
  • 75485 10/31/01 10:49 Jack Bubbler conspiracy
  • 75481 10/31/01 09:42 LeRose Left Dipole NMR
  • 75476 10/31/01 08:06 reitz owl shift summary
  • 75474 10/31/01 05:50 Reitz Left Dipole
  • 75471 10/31/01 00:05 Adaq 10/30 day shift summary
  • Tuesday

  • 75469 10/30/01 21:17 Calarco End of run entries
  • 75468 10/30/01 20:16 johna Clarification on HRS angles.
  • 75467 10/30/01 18:32 Jack Left arm VDC power supply
  • 75466 10/30/01 18:12 johna Problem in injector, probably down at least an hour.
  • 75465 10/30/01 18:11 johna HRS angles from camera.
  • 75464 10/30/01 17:52 johna Getting ready for beam
  • 75462 10/30/01 14:54 Xiaofeng Z Owl shift summary
  • 75460 10/30/01 13:18 Reitz RICH work
  • 75448 10/30/01 12:49 R. Michaels R-arm Portserver ok again
  • 75447 10/30/01 11:21 jpchen analyzed runs 1339-1349
  • 75445 10/30/01 07:13 Xiaofeng Zhu Beam is off at 7:00am
  • 75443 10/30/01 06:46 Xiaofeng OTR radiator masked for beam
  • 75441 10/30/01 05:50 Xiaofeng Network problem
  • 75437 10/30/01 04:48 Xiaofeng Beam tuning for 10 minutes
  • 75428 10/30/01 00:19 weinstein run comment wrong
  • 75425 10/30/01 00:02 weinstein shift summary
  • Monday

  • 75407 10/29/01 17:12 weinstein detectors
  • 75406 10/29/01 17:05 weinstein VDCX
  • 75402 10/29/01 16:20 Saha Shift Summary
  • 75401 10/29/01 16:10 Nilanga Beta optimization
  • 75400 10/29/01 16:06 Nilanga coincidence TOF
  • 75389 10/29/01 12:17 Saha Arc Integral Measurements
  • 75386 10/29/01 11:41 Nilanga Emiss analysis
  • 75383 10/29/01 10:00 S. Sirca OTR values in BPM/HARP/OTR scans
  • 75367 10/29/01 04:46 Hyde-Wright magnets
  • 75366 10/29/01 04:30 Hyde-Wright magnets
  • 75357 10/29/01 01:43 Hyde-Wright OTR
  • 75350 10/29/01 00:06 x. jiang shift summary
  • Sunday

  • 75342 10/28/01 21:39 Wendy Beta optimization
  • 75340 10/28/01 20:59 x. jiang beam will be gone for one hour.
  • 75321 10/28/01 15:57 Calarco Shift summary
  • 75317 10/28/01 14:13 Calarco Q2-left trip
  • 75315 10/28/01 13:37 Calarco Spectrometer angle readout
  • 75313 10/28/01 13:13 Calarco Moeller target
  • 75312 10/28/01 12:50 Calarco Kinematics change, Q2 trip
  • 75309 10/28/01 11:58 Calarco Magnets recovered
  • 75303 10/28/01 10:30 S. Sirca Cryo problems due to communication failures solved
  • 75297 10/28/01 09:31 Calarco Spectrometer cryogen levels
  • 75296 10/28/01 09:01 Calarco HRS magnet status
  • 75295 10/28/01 08:22 R. Michaels DAQ problems, fixed
  • 75291 10/28/01 07:33 Mattias Shift summary
  • 75290 10/28/01 07:13 S. Sirca IOC/network problems
  • 75271 10/28/01 02:58 Mattias Physics production
  • 75268 10/28/01 02:42 Mattias OTR 1C12 out
  • 75263 10/28/01 02:11 Mattias Winter time change
  • 75262 10/28/01 01:58 Mattias Target thickness calibration finished
  • 75250 10/28/01 00:02 x. jiang shift summary
  • 75249 10/28/01 00:02 reitz BeO
  • 75248 10/28/01 00:00 S. Sirca Right HRS floor-marks
  • Saturday

  • 75243 10/27/01 22:56 Reitz hall a main screen at 59.97 deg
  • 75239 10/27/01 22:31 x. jiang 1301 is a junk run
  • 75234 10/27/01 22:23 x. jiang tripple images in phi_t
  • 75232 10/27/01 21:39 x. jiang double images in phi_t.
  • 75230 10/27/01 20:08 x. jiang run type are optics run 1291-1298
  • 75221 10/27/01 19:06 x. jiang run 1292, image of collimator
  • 75213 10/27/01 18:21 Reitz magnet screen for optics run
  • 75212 10/27/01 17:56 X. Jiang left arm sieve slit pattern in run 1291
  • 75207 10/27/01 17:18 Reitz Magnet screen for RCS optics
  • 75205 10/27/01 16:07 bogdanw shift summary
  • 75203 10/27/01 15:29 Calarco Q1-Left Arm problem
  • 75202 10/27/01 14:30 Calarco Q1-Left Arm
  • 75175 10/27/01 08:01 R. Michaels owl shift summary
  • 75152 10/27/01 05:10 Mattias Using lower current
  • 75150 10/27/01 04:37 R. Michaels White spectra R done, going to L
  • 75140 10/27/01 03:49 Adaq correction
  • 75137 10/27/01 03:34 R. Michaels results of target current scan
  • 75128 10/27/01 02:45 Mattias Shower counter set values for HV changed, but still off.
  • 75123 10/27/01 01:57 R. Michaels More accurate Pump Speed Optimization
  • 75122 10/27/01 01:49 Mattias Correction
  • 75112 10/27/01 00:46 Mattias Start of Run Halog momentum entries wrong!
  • 75107 10/27/01 00:04 Reitz swing shift summary
  • 75106 10/27/01 00:01 reitz first results on beam current scan
  • Friday

  • 75102 10/26/01 23:48 reitz magnets for 1.098 / 4.12
  • 75100 10/26/01 23:31 reitz run 1242 prescaling
  • 75094 10/26/01 22:39 Moussiegt magnet screen for white scan
  • 75093 10/26/01 21:13 Reitz maximum beam 80 uA
  • 75092 10/26/01 20:56 Reitz right arm Q2 quenched
  • 75078 10/26/01 17:28 atrig Trigger-Download
  • 75076 10/26/01 15:31 bogdanw shift summary
  • 75074 10/26/01 14:43 Eugene Chuda Shower detector (right arm) turned off
  • 75073 10/26/01 13:59 Calarco Beam availability
  • 75072 10/26/01 12:06 S. Sirca BPM values in BPM/HARP scans
  • 75061 10/26/01 07:59 R. Michaels owl shift summary
  • 75056 10/26/01 07:33 Mattias PUMPSET = 1200
  • 75054 10/26/01 07:17 R. Michaels pump speed (cont) and a graph
  • 75052 10/26/01 07:07 Mattias Pump Speed Cal 1750
  • 75048 10/26/01 06:54 Mattias Pump Speed Cal 1500
  • 75046 10/26/01 06:39 Mattias Pump speed cal 1250
  • 75042 10/26/01 06:30 Mattias Pump speed cal 1000
  • 75039 10/26/01 05:32 R. Michaels more tgt pump speed calib. Then no beam.
  • 75035 10/26/01 04:34 Mattias Pump speed Cal 750
  • 75033 10/26/01 04:25 R. Michaels Target pump speed calib. Aborted, no beam
  • 75030 10/26/01 03:02 Mattias Pumpspeed Cal 500
  • 75026 10/26/01 02:43 R. Michaels no BCM calib, do pump speed scan & re: Ytgt bgr.
  • 75025 10/26/01 02:36 Wendy Beam Centering Study
  • 75022 10/26/01 02:02 R. Michaels 0,0 and no water
  • 75020 10/26/01 02:01 R. Michaels post scan -1.4, 0. & some conclusions
  • 75016 10/26/01 01:38 R. Michaels Post scan
  • 75014 10/26/01 01:32 R. Michaels Post scan, +2, 0
  • 75010 10/26/01 01:15 R. Michaels Post scan +1,0
  • 75008 10/26/01 00:57 R. Michaels Post scan, step 0
  • 75007 10/26/01 00:50 R. Michaels here is BPM fig
  • 75006 10/26/01 00:48 R. Michaels HARP scan
  • 75002 10/26/01 00:01 Rikki shift summary
  • Thursday

  • 74999 10/25/01 23:30 Rikki HARP - beam pos -1,0
  • 74996 10/25/01 23:23 Rikki HARP - beam position 1,0
  • 74994 10/25/01 23:01 Rikki HARP scan - beam pos 0,-1
  • 74990 10/25/01 22:56 Rikki harp scan - beam pos. 0,1
  • 74987 10/25/01 22:47 Rikki harp scan - beam position 0,0
  • 74984 10/25/01 20:27 Rikki Harp scan #1412
  • 74979 10/25/01 17:01 Nilanga U1 plane Software problem fixed.
  • 74978 10/25/01 16:40 Higinbotham Harp #2 Fixed
  • 74977 10/25/01 16:32 atrig Trigger-Download
  • 74973 10/25/01 16:19 weinstei shift summary
  • 74971 10/25/01 12:07 S. Sirca Right Q1 cryo problem fixed
  • 74967 10/25/01 08:06 Mattias Shift Summary
  • 74966 10/25/01 02:03 Mattias Waterfall Target Screen Dump
  • 74965 10/25/01 01:56 Mattias Left Arm HV Screen Dump
  • 74964 10/25/01 01:55 Mattias Right Arm HV Screen Dumps
  • 74963 10/25/01 01:45 Mattias ARC Field Screen Dump
  • 74962 10/25/01 01:42 Mattias ARC Scan Screen Dumps
  • 74961 10/25/01 01:35 Mattias Screen dump of General
  • 74960 10/25/01 01:18 Higinbotham Q1 Problem Continues
  • 74959 10/25/01 01:10 Higinbotham Harp #2 Not Working
  • 74958 10/25/01 00:30 Higinbotham Q1 Tripped
  • Wednesday

  • 74956 10/24/01 23:44 Mattias Shower counter HV increased, new values
  • 74955 10/24/01 23:32 Mattias Shower counter voltages increased
  • 74953 10/24/01 18:17 weinstei magnet cryogen
  • 74952 10/24/01 18:09 weinstei magnet coolant
  • 74951 10/24/01 17:18 Mattias Magnet N2 level low
  • 74950 10/24/01 16:59 Mattias Halla A Lock up
  • 74948 10/24/01 15:36 Reitz Trigger Downloads
  • 74946 10/24/01 15:31 atrig Trigger-Download
  • 74945 10/24/01 15:29 Mattias S0 HV turned on
  • 74943 10/24/01 11:22 atrig Trigger-Download
  • Tuesday

  • 74941 10/23/01 22:29 Mattias Shower Counter HV turned on
  • 74939 10/23/01 21:55 aregjan left HRS changed to 1GeV
  • 74938 10/23/01 20:58 bogdanw left arm momentum
  • 74936 10/23/01 18:50 fissum S0 HVs turned off
  • Monday

  • 74912 10/22/01 16:25 Mattias Ande Test
  • 74909 10/22/01 13:00 Reitz test of withrich configuration
  • Sunday

  • 74905 10/21/01 15:57 R. Michaels BPM readout modification
  • Saturday

    
    

    Friday

  • 74869 10/19/01 18:57 atrig Trigger-Download
  • 74868 10/19/01 18:29 atrig Trigger-Download
  • Thursday

  • 74865 10/18/01 14:59 atrig Trigger-Download
  • Wednesday

    
    

    Tuesday

  • 74827 10/16/01 11:59 LeRose Left Arm max P0
  • Monday

  • 74817 10/15/01 21:24 atrig Trigger-Download
  • 74805 10/15/01 12:45 LeRose Left dipole controls
  • 74804 10/15/01 12:42 LeRose Left Arm High P
  • Saturday

    
    

    Friday

    
    

    Thursday

    
    

    Wednesday

  • 74768 10/10/01 15:19 R. Michaels/ trigger checkout
  • 74765 10/10/01 14:36 atrig Trigger-Download
  • 74762 10/10/01 14:32 atrig Trigger-Download
  • 74760 10/10/01 14:28 atrig Trigger-Download
  • 74757 10/10/01 14:22 atrig Trigger-Download
  • 74753 10/10/01 14:17 atrig Trigger-Download
  • 74749 10/10/01 14:10 atrig Trigger-Download
  • 74746 10/10/01 14:07 atrig Trigger-Download
  • Tuesday

    
    

    Monday

    
    

    Thursday

  • 74676 10/04/01 19:20 apar Start_of_Run_Parity_3613
  • 74675 10/04/01 19:11 apar Start_of_Run_Parity_3612
  • 74674 10/04/01 19:02 apar Start_of_Run_Parity_3611
  • 74673 10/04/01 18:54 apar Start_of_Run_Parity_3610
  • 74672 10/04/01 18:42 apar Start_of_Run_Parity_3609
  • 74671 10/04/01 18:26 apar Start_of_Run_Parity_3608
  • 74670 10/04/01 18:17 apar Start_of_Run_Parity_3607
  • 74669 10/04/01 18:05 apar Start_of_Run_Parity_3606
  • 74668 10/04/01 17:55 apar Start_of_Run_Parity_3605
  • 74667 10/04/01 17:45 apar Start_of_Run_Parity_3604
  • 74666 10/04/01 17:25 apar Start_of_Run_Parity_3603
  • 74665 10/04/01 17:14 apar Start_of_Run_Parity_3602
  • 74664 10/04/01 17:02 apar Start_of_Run_Parity_3601
  • 74663 10/04/01 16:49 apar Start_of_Run_Parity_3600
  • 74662 10/04/01 16:37 apar Start_of_Run_Parity_3599
  • 74661 10/04/01 16:26 apar Start_of_Run_Parity_3598
  • 74660 10/04/01 16:18 apar Start_of_Run_Parity_3597
  • 74659 10/04/01 16:15 apar Start_of_Run_Parity_3596
  • 74658 10/04/01 16:01 apar Start_of_Run_Parity_3595
  • 74657 10/04/01 15:55 apar Start_of_Run_Parity_3594
  • 74656 10/04/01 15:28 apar Start_of_Run_Parity_3593
  • Wednesday

  • 74632 10/03/01 11:30 atrig Trigger-Download
  • Tuesday

    
    

    Monday

  • 74627 10/01/01 13:03 atrig Trigger-Download